# One-Step Model Predictive Path Integral for Manipulator Motion Planning Using Configuration Space Distance Fields

### v1

@@ Abstract
Motion planning for robotic manipulators is a fundamental problem in robotics. Classical optimization-based methods typically rely on the gradients of signed distance fields (SDFs) to impose collision-avoidance constraints. However, these methods are susceptible to local minima and may fail when the SDF gradients vanish. Recently, Configuration Space Distance Fields (CDFs) have been introduced, which directly model distances in the robot’s configuration space. Unlike workspace SDFs, CDFs are differentiable almost everywhere and thus provide reliable gradient information. On the other hand, gradient-free approaches such as Model Predictive Path Integral (MPPI) control leverage long-horizon rollouts to achieve collision avoidance. While effective, these methods are computationally expensive due to the large number of trajectory samples, repeated collision checks, and the difficulty of designing cost functions with heterogeneous physical units. In this paper, we propose a framework that integrates CDFs with MPPI to enable direct navigation in the robot’s configuration space. Leveraging CDF gradients, we unify the MPPI cost in joint-space and reduce the horizon to one step, substantially cutting computation while preserving collision avoidance in practice. We demonstrate that our approach achieves nearly 100% success rates in 2D environments and consistently high success rates in challenging 7-DOF Franka manipulator simulations with complex obstacles. Furthermore, our method attains control frequencies exceeding 750 Hz, substantially outperforming both optimization-based and standard MPPI baselines. These results highlight the effectiveness and efficiency of the proposed CDF-MPPI framework for high-dimensional motion planning.
@@ I Introduction
Motion planning is one of the most active research areas in robotic control, with extensive efforts devoted to achieving efficient and collision-free robot motions. Many studies [1, 2] formulate this problem as an optimization problem, where the signed distance field (SDF) is used to represent obstacle proximity and its gradient is employed to guide the optimization toward collision-free trajectories. In practice, classical approaches such as CHOMP [2] precompute SDFs on a voxel grid via the Euclidean Distance Transform (EDT), while TrajOpt [1] repeatedly queries distances using GJK [3] between convex bodies. Both computations are typically performed on CPUs, which is sufficient for static environments and moderate problem sizes.
More recent works, such as [4], which adopt regression-based formulations, and [5], which leverage neural networks, demonstrate that SDFs can be efficiently fitted and evaluated on GPUs. These approaches enable faster and more accurate SDF computation. However, as in previous studies, gradient information is still required to achieve collision-free motion. For robotic arms, this introduces two issues: (i) the SDF gradient may vanish near obstacles [6], and (ii) the distribution of SDF values in the joint space is highly non-uniform [7].
Figure 1: Comparison between the original MPPI and our proposed method in the configuration space. The irregular shapes on the left represent obstacles, blue thin lines denote sampled trajectories at each step, the black line indicates the resulting trajectory, and the white dot marks the start configuration. Left: the original MPPI samples long-horizon trajectories at each step. Right: our method leverages CDFs to sample one-step horizon trajectories at each step.
To overcome these limitations, sampling-based methods have emerged as a viable alternative. Their key advantage lies in the fact that they do not require gradient information, thereby avoiding issues such as the vanishing-gradient problem in SDFs. Representative approaches include Rapidly-exploring Random Trees (RRT) [8] and Probabilistic Roadmaps (PRM) [9]. However, these methods are often inefficient due to their reliance on extensive random sampling.
@@
Configuration Space Distance Fields (CDFs), proposed in [7], provide a promising alternative. Unlike traditional SDFs defined in the workspace, CDFs directly construct distance fields in the configuration space, thereby addressing both the non-uniformity and vanishing-gradient problems. CDFs thus offer an attractive framework for robot motion planning, as their gradients can be directly exploited for navigation in configuration space [7]. However, optimization-based methods that rely on CDFs are susceptible to local minima, which may result in navigation failures. Furthermore, such optimization is typically executed on CPUs, making it time-consuming.
In this work, we integrate the MPPI framework with CDFs and propose a simple yet efficient algorithm that addresses several limitations of existing methods:
@@
    Unified and simplified cost function: Leveraging the property that CDFs provide well-distributed, non-vanishing gradients in the configuration space, we redesign the MPPI cost function. In particular, we employ angular navigation to normalize and unify the scale of different cost terms.
@@
    Faster computation: The redesigned cost function enables one-step trajectory sampling instead of requiring long-horizon rollouts, thereby significantly reducing computation time. Most computations can be efficiently executed on GPUs (Fig. 1 illustrates a comparison between our method and the original MPPI).
@@ II-A Model Predictive Path Integral
The MPPI framework operates as follows. For a discrete-time robotic system at time step $t$, the control input $u$ represents the joint velocity. The input $u$ is sampled from a policy $\pi_{\phi_{t}}=\prod_{h=0}^{H-1}\pi_{\phi_{t},h}$ where $H$ is the prediction horizon and $\pi_{\phi_{t},h}$ is a Gaussian distribution with mean $\mu_{\phi_{t},h}$ and covariance $\Sigma_{\phi_{t},h}$.
At each time step $t$, the sampling-based MPC algorithm proposed by [13] draws a batch of $N$ control sequences ${u}_{i=1..N}^{h=1..H}$ from the current policy $\pi_{\phi_{t}}$. For each sampled sequence, the corresponding cost $c_{i,h}$} is computed from the state trajectories ${x}_{i,h}$ obtained by propagating the dynamics of the robot using the sampled controls.
The Gaussian policy parameters $\pi_{\phi_{t},h}=\mathcal{N}(\mu_{t,h},\Sigma_{t,h})$ are then updated according to the following weighted rules:
$\mu_{t+1,h}\leftarrow$ $(1-\alpha_{\mu})\mu_{t,h}+\alpha_{\mu}\frac{\sum_{i=1}^{N}w_{i}u_{i,h}}{\sum_{i=1}^{N}w_{i}},$ (1a)
$\Sigma_{t+1,h}\leftarrow$ $(1-\alpha_{\Sigma})\Sigma_{t,h}$
$+\alpha_{\Sigma}\frac{\sum_{i=1}^{N}w_{i}(u_{i,h}-\mu_{t,h})(u_{i,h}-\mu_{t,h})^{\top}}{\sum_{i=1}^{N}w_{i}}.$ (1b)
where $\alpha_{\mu}$ and $\alpha_{\Sigma}$ are filtering coefficients. The weight for each sample is defined as $w_{i}=\exp\left(-\frac{1}{\beta}\hat{C}_{i}\right)$, where $\beta$ is a temperature parameter for the cost. The cost value $\hat{C}(x_{t},u)$ is computed as
$\hat{C}(x_{t},u)=\sum_{h=0}^{H-2}\gamma^{h}\hat{c}(x_{n,h},u_{n,h})+\gamma^{H-1}\hat{c}(x_{n,H-1}),$ (2)
@@
The stage cost $c_{x,u}$ is composed of multiple terms, including joint-limit avoidance, self-collision and collision avoidance, contingency stopping, and manipulation cost:
$c_{x,u}=\alpha_{s}c_{stop}+\alpha_{m}c_{mani}+\alpha_{c}c_{col}+\alpha_{j}c_{joint}.$ (3)
@@ II-B Configuration Space Distance Fields
subject to the constraint that $f_{s}(\mathbf{p},\mathbf{q}^{\prime})=0$, where $\mathbf{q}^{\prime}$ denotes a configuration in contact with the obstacle and $\mathbf{p}$ is a corresponding contact point in the workspace. In practice, $f_{c}(\mathbf{p},\mathbf{q})$ can be encoded using a neural network. The CDF possesses several desirable properties. In particular, it satisfies the eikonal equation $\|\nabla_{\mathbf{q}}f_{c}(\mathbf{p},\mathbf{q})\|=1$ almost everywhere in the configuration space. This property makes CDFs especially well-suited for motion planning, since the gradient $\nabla_{\mathbf{q}}f_{c}$ naturally points in the locally optimal escape direction away from the nearest obstacle, thereby providing a direct guidance signal for collision avoidance.
@@ III Method
We aim to control an $n$-DOF robotic arm to perform fast, collision-free motions. In this section, we present a method that combines MPPI and CDFs to achieve this goal.
@@
where $\mathbf{q}_{t}$ is the joint configuration at time step $t$, $\mathbf{u}$ is the joint input, and $\Delta t$ is the discrete control time step.
@@ III-A Improved Model Predictive Path Integral with an Angle-Based Cost Function
We propose an angle-based cost function that unifies the cost terms into a common unit. Let $\mathbf{q}_{t}$ be the robot configuration at time $t$, $\mathbf{q}_{t+1}$ the configuration at the next time step, $\mathbf{q}_{f}$ the goal configuration, and $\mathbf{p}$ is the obstacle point in the work space. The proposed cost is defined as:
@@
where $\alpha_{1},\alpha_{2}>0$ are scalar weights. $\theta_{1}\in[0,\pi]$ is the angle between the obstacle normal direction at $\mathbf{q}_{t}$ and the motion vector $\mathbf{q}_{next}=(\mathbf{q}_{t+1}-\mathbf{q}_{t})$, while $\theta_{2}\in[0,\pi]$ is the angle between the goal vector $\mathbf{e}_{t}=(\mathbf{q}_{f}-\mathbf{q}_{t})$ and the motion vector $(\mathbf{q}_{t+1}-\mathbf{q}_{t})$. When $\theta_{2}=0$, the robot moves directly toward the goal; $\theta_{2}=\pi$ indicates motion directly away from it. And $\theta_{2}$ can be computed as:
$\theta_{2}(\mathbf{q}_{t},\mathbf{q}_{t+1})$ $=\arccos\frac{\mathbf{q}_{next}\cdot\mathbf{e}_{t}}{\|\mathbf{q}_{next}\|\,\|\mathbf{e}_{t}\|}$ (8)
@@
  • Goal proximity: If the distance to the obstacle is larger than the distance to the goal, i.e., $f_{c}(\mathcal{P},\mathbf{q}_{t})\geq\lVert\mathbf{q}_{goal}-\mathbf{q}_{t}\rVert$, as illustrated in Fig. 2(Right), it is more efficient to move directly towards the goal rather than proactively avoiding the obstacle.
Figure 2: Left: The blue arrow denotes the actual motion direction of the robot in the configuration space at the next step, forming two angles with the CDF gradient and the goal state. Right: Illustration of why the obstacle-avoidance angle cost is set to zero when the robot–obstacle distance exceeds the robot–goal distance.
This design offers two advantages: (1) all cost terms are expressed in the same unit (angle), making the weight selection more intuitive; (2) obstacle avoidance and goal-seeking behaviors are directly captured in a single unified metric. In MPPI, a batch of $N$ control sequences is sampled from $\pi_{\phi_{t}}$, and each sequence is rolled out over a prediction horizon $H$ to compute the total cost. With the proposed angle-based cost, collision risk and goal progress can be evaluated locally at the current time step, without simulating the full horizon. Therefore, we modify the sampling policy to generate only $N$ single-step control inputs ($H=1$) from $\pi_{\phi_{t}}$, compute $\mathbf{q}_{t+1}-\mathbf{q}_{t}$, and evaluate the cost directly. This reduces computation while retaining effective obstacle avoidance and goal-directed motion.
@@ III-B One-step MPPI for Configuration Space Motion Planning
The projection of any control $\mathbf{u}\in\mathbb{R}^{m}$ onto $\mathcal{U}_{\text{step}}(\mathbf{q})$ is the elementwise clipping
@@ IV Numerical Simulations
In this section, we report simulation results on a two-link robot and a Franka Emika robot to evaluate the proposed method. The experiments are designed to (i) demonstrate the necessity of updating both the mean and covariance of the sampling distribution, (ii) show that naïve sampling without distribution updates is insufficient, (iii) assess the effectiveness of our cost-function design, and (iv) compare our approach against representative baselines. Finally, we provide a qualitative analysis to further validate the method. All experiments were performed on a workstation with an AMD Ryzen 700 CPU and an NVIDIA RTX 4070 GPU.
@@
0: Initial configuration $\mathbf{q}_{\mathrm{init}}$; goal configuration $\mathbf{q}_{\mathrm{goal}}$; obstacle point cloud $\mathcal{P}$; joint limits $(\mathbf{q}_{\min},\mathbf{q}_{\max})$; time step $\Delta t$; number of samples NN; angle-based cost weights $(\alpha_{1},\alpha_{2})$; obstacle threshold distance $d_{act}$. MPPI parameters.
@@
1: Initialize $\mathbf{q}_{0}\leftarrow\mathbf{q}_{\mathrm{init}}$
2: Initialize Gaussian policy $\pi_{\phi_{0}}=\mathcal{N}(\mu_{0},\Sigma_{0})$
@@
4:  Sample NN control inputs $\{\mathbf{u}_{i}\}_{i=1}^{N}\sim\pi_{\phi_{t}}$ {one-step (H=1H=1)}
@@
11:  Update policy parameters $\mu_{t+1}$, $\Sigma_{t+1}$ using Eq. 1a and Eq. 1b
12:  Set the next control input $\mathbf{u}_{\mathrm{des}}\leftarrow\mu_{t+1}$
@@
14:  if $\lVert\mathbf{q}_{t+1}-\mathbf{q}_{\mathrm{goal}}\rVert<allow\_range$ then
@@
IV-A 2D Two-Link Robot
We first evaluate the proposed method on a 2D two-link robot with link lengths $l_{1}=l_{2}=2$ and joint limits $q\in[-\pi,\pi]$ for each joint. Two circular obstacles are placed at positions (2.3,2.3)(2.3,-2.3) and $(0.0,2.45)$, each with a radius of $0.3$. The start configuration is $\mathbf{q}_{\mathrm{start}}=[2.1,1.2]$, and the goal configurations are $\mathbf{q}_{\mathrm{goal}}=[-2.1,-0.9]$ and [0.5,0.0][-0.5,0.0]. We set the control time step to $\Delta t=0.01$
@@
We further provide a comprehensive comparison with other baseline methods: the original MPPI [13] and the QP-based method from [7], solved using IPOPT [14]. The start and goal configurations are the same as in the previous experiment, with an additional test case for the QP method (labeled QP 1), where the start is $[2.5,0.5]$ and the goal is [2.0,0.3][-2.0,0.3]. The cost function used in the MPPI experiment is defined as
$c_{x,u}=\alpha_{g}c_{\text{goal}}+\alpha_{c}c_{\text{col}}+\alpha_{j}c_{\text{joint}}+\alpha_{g}c_{\text{stay}}.$ (16)
where $c_{\text{goal}}$ denotes the terminal cost measuring the error between the predicted final configuration and the target configuration $\mathbf{q}_{\mathrm{goal}}$, while $c_{\text{stay}}$ penalizes trajectories that remain stationary over time [4].
The trajectories in configuration space are shown in Fig. 3(Right). The QP-based method [7] occasionally fails to reach the goal for certain start–goal pairs, as the control input oscillates between two states, leading to stagnation and convergence to a local optimum (QP 2, same start and goal as MPPI (Ours) and MPPI (Original)). The original MPPI produces trajectories that often run close to obstacles and joint limits, and the resulting paths are generally less smooth.
Figure 3: Comparison of different methods in both configuration space and workspace trajectories. Left: Trajectories from the same start configuration but different goals, used to validate the effectiveness of our cost function design. A purely random baseline, which samples controls independently at each step without policy updates, highlights the necessity of distribution updates in MPPI. Right: Comparison with MPPI and an optimization-based method. For QP2 (black), the start and goal are identical to those used in MPPI (Original) and MPPI (Ours), but the trajectory converges to a local optimum and fails to progress. In contrast, QP1, with different start and goal states, successfully reaches the target.
@@
measured in radians in the joint space. All methods are tested under the same random seed with identical sampling rules for start and goal states, repeated 500 times. The results are summarized in Table. I. Our proposed method achieves the highest success rate. Although it requires more steps and produces slightly longer paths, this can be attributed to its ability to avoid local optima and handle longer trajectories effectively. The failure cases mainly occur when the start configuration is very close to an obstacle, causing collisions before the distribution has been sufficiently updated. In contrast, the original MPPI may still result in collisions with obstacles or violations of joint limits. A detailed runtime analysis is presented in the following subsection, demonstrating that our method maintains computational efficiency.
@@
$\mathbf{u}=k(\mathbf{q}_{\text{goal}}-\mathbf{q}_{\text{start}}).$ (18)

### v2

@@ Abstract
Motion planning for robotic manipulators is a fundamental problem in robotics. Classical optimization-based methods typically rely on the gradients of signed distance fields (SDF) to impose collision-avoidance constraints. However, these methods are susceptible to local minima and may fail when the SDF gradients vanish. Recently, Configuration Space Distance Fields (CDF) have been introduced, which directly model distances in the robot’s configuration space. Unlike workspace SDF, CDF are differentiable almost everywhere and thus provide reliable gradient information. On the other hand, gradient-free approaches such as Model Predictive Path Integral (MPPI) control leverage long-horizon rollouts to achieve collision avoidance. While effective, these methods are computationally expensive due to the large number of trajectory samples, repeated collision checks, and the difficulty of designing cost functions with heterogeneous physical units. In this paper, we propose a framework that integrates CDF with MPPI to enable direct navigation in the robot’s configuration space. Leveraging CDF gradients, we unify the MPPI cost in joint-space and reduce the horizon to one step, substantially cutting computation while preserving collision avoidance in practice. We demonstrate that our approach achieves nearly 100% success rates in 2D environments and consistently high success rates in challenging 7-DOF Franka manipulator simulations with complex obstacles. Furthermore, our method attains control frequencies exceeding 750 Hz, substantially outperforming both optimization-based and standard MPPI baselines. These results highlight the effectiveness and efficiency of the proposed CDF-MPPI framework for high-dimensional motion planning.
@@ I Introduction
Motion planning is one of the most active research areas in robotic control, with extensive efforts devoted to achieving efficient and collision-free robot motions. Many studies [1, 2] formulate this problem as an optimization problem, where the signed distance field (SDF) is used to represent obstacle proximity and its gradient is employed to guide the optimization toward collision-free trajectories. In practice, classical approaches such as CHOMP [2] precompute SDF on a voxel grid via the Euclidean Distance Transform (EDT), while TrajOpt [1] repeatedly queries distances using GJK [3] between convex bodies. These computations are often implemented on CPUs in standard libraries, which is generally sufficient for static environments and moderate-scale problems. Recent work has explored GPU acceleration of both SDF construction and GJK queries to handle dynamic scenes and large-scale planning more efficiently, though they can still be computationally expensive due to frequent voxel updates and the iterative nature of distance queries.
More recent works, such as [4], which adopt regression-based formulations, and [5], which leverage neural networks, demonstrate that SDF can be efficiently fitted and evaluated on GPU. These approaches enable faster and more accurate SDF computation. However, as in previous studies, gradient information is required to achieve collision-free motion. For robotic arms, this introduces two issues: (i) the SDF gradient may vanish near obstacles [6], and (ii) the distribution of SDF values in the configuration space is highly non-uniform [7].
Figure 1: Comparison between the original MPPI and our proposed method in the two-link robot configuration space. The irregular shapes on the left represent obstacles, blue thin lines denote sampled trajectories at each step, the black line indicates the resulting trajectory, and the white dot marks the start configuration. Left: the original MPPI samples long-horizon trajectories at each step. Right: our method leverages CDF to sample one-step horizon trajectories at each step.
To overcome these limitations, sampling-based methods have emerged as a viable alternative. Their key advantage lies in the fact that they do not require gradient information, thereby avoiding issues such as the vanishing-gradient problem in SDF. Representative approaches include Rapidly-exploring Random Trees (RRT) [8] and Probabilistic Roadmaps (PRM) [9]. However, these methods are often inefficient due to their reliance on extensive random sampling.
@@
Configuration Space Distance Fields (CDF), proposed in [7], provide a promising alternative. Unlike traditional SDF defined in the workspace, CDF directly construct distance fields in the configuration space, thereby addressing both the non-uniformity and vanishing-gradient problems. CDF thus offer an attractive framework for robot motion planning, as their gradients can be directly exploited for navigation in configuration space [7]. However, optimization-based methods that rely on CDF are susceptible to local minima, which may result in navigation failures. Moreover, such optimization problems typically require repeated iterations, making the computation slow especially in high-dimensional configuration spaces.
In this work, we integrate the MPPI framework with CDF and propose a simple yet efficient algorithm that addresses several limitations of existing methods. Our goal is to control an $n$-DOF robotic arm to perform fast and collision-free motions.
@@
    Unified and simplified cost function: Leveraging the property that CDF provide well-distributed, non-vanishing gradients in the configuration space, we redesign the MPPI cost function. In particular, we employ angular navigation to normalize and unify the scale of different cost terms.
@@
    Faster computation: The redesigned cost function enables one-step trajectory sampling instead of requiring long-horizon rollouts, thereby significantly reducing computation time. Most computations can be efficiently executed on GPU (Fig. 1 illustrates a comparison between our method and the original MPPI).
@@ II-A Model Predictive Path Integral
The MPPI framework operates as follows. For a discrete-time robotic system at time step $t$, the control input $\mathbf{u}$} represents the joint velocity. The input $\mathbf{u}$} is sampled from a policy $\pi_{t}=\prod_{h=0}^{H-1}\pi_{t,h}$ where $H$ is the prediction horizon and $\pi_{t,h}$ is a Gaussian distribution with mean $\boldsymbol{\mu}_{t,h}$ and covariance $\boldsymbol{{\boldsymbol{\Sigma}}}_{t,h}$.
At each time step $t$, the MPPI algorithm proposed by [13] draws a batch of $N$ control sequences ${{\mathbf{u}}}_{i=1..N}^{h=1..H}$ from the current policy $\pi_{t}$. For each sampled sequence, the corresponding cost $c_{i,h}$} is computed from the state trajectories ${x}_{i,h}$ obtained by propagating the dynamics of the robot using the sampled controls.
The Gaussian policy parameters $\pi_{t,h}=\mathcal{N}(\boldsymbol{\mu_{t,h}},{\boldsymbol{\Sigma}}_{t,h})$ are then updated according to the following weighted rules:
$\boldsymbol{\mu}_{t+1,h}\leftarrow$ $(1-\alpha_{\mu}){\boldsymbol{\mu}}_{t,h}+\alpha_{\boldsymbol{\mu}}\frac{\sum_{i=1}^{N}w_{i}\mathbf{u}_{i,h}}{\sum_{i=1}^{N}w_{i}},$ (1a)
${\boldsymbol{\Sigma}}_{t+1,h}\leftarrow$ $(1-\alpha_{\boldsymbol{\Sigma}}){\boldsymbol{\Sigma}}_{t,h}$
$+\alpha_{\boldsymbol{\Sigma}}\frac{\sum_{i=1}^{N}w_{i}(\mathbf{u}_{i,h}-{\boldsymbol{\mu}}_{t,h})(\mathbf{u}_{i,h}-{\boldsymbol{\mu}}_{t,h})^{\top}}{\sum_{i=1}^{N}w_{i}}.$ (1b)
where $\alpha_{\boldsymbol{\mu}}$ and $\alpha_{\boldsymbol{\Sigma}}$ are filtering coefficients. The weight for each sample is defined as $w_{i}=\exp\left(-\frac{1}{\beta}\hat{C}_{i}\right)$, where $\beta$ is a temperature parameter for the cost. The cost value $\hat{C}(\mathbf{x}_{t},{\mathbf{u}})$ is computed as
$\hat{C}(\mathbf{x}_{t},{\mathbf{u}})=\sum_{h=0}^{H-2}\gamma^{h}c(\mathbf{x}_{n,h},\mathbf{u}_{n,h})+\gamma^{H-1}\hat{c}(\mathbf{x}_{n,H-1},\mathbf{u}_{n,H-1}),$ (2)
@@
The stage cost $c_{\mathbf{x},{\mathbf{u}}}$ is composed of multiple terms, including joint-limit avoidance, self-collision and collision avoidance, contingency stopping, and manipulation cost:
$c_{\mathbf{x},{\mathbf{u}}}=\alpha_{s}c_{stop}+\alpha_{m}c_{mani}+\alpha_{c}c_{col}+\alpha_{j}c_{joint}.$ (3)
@@ II-B Configuration Space Distance Fields
subject to the constraint that $f_{s}(\mathbf{p},\mathbf{q}^{\prime})=0$, where $\mathbf{q}^{\prime}$ denotes a configuration in contact with the obstacle and $\mathbf{p}$ is a corresponding contact point in the workspace. In practice, $f_{c}(\mathbf{p},\mathbf{q})$ can be encoded using a neural network. The CDF possesses several desirable properties. In particular, it satisfies the eikonal equation $\|\nabla_{\mathbf{q}}f_{c}(\mathbf{p},\mathbf{q})\|=1$ almost everywhere in the configuration space. This property makes CDF especially well-suited for motion planning, since the gradient $\nabla_{\mathbf{q}}f_{c}$ naturally points in the locally optimal escape direction away from the nearest obstacle, thereby providing a direct guidance signal for collision avoidance.
@@ III Method
We aim to control an $n$-DOF robotic arm to perform fast, collision-free motions. In this section, we present a method that combines MPPI and CDF to achieve this goal.
@@
where $\mathbf{q}_{t}$ be the robot configuration at time $t$, $\mathbf{q}_{t+1}$ the configuration at the next time step, $\mathbf{u}$ is the joint input, and $\Delta t$ is the discrete control time step.
@@ III-A Improved Model Predictive Path Integral with an Angle-Based Cost Function
We propose an angle-based cost function that unifies the cost terms into a common unit. Let $\mathbf{q}_{f}$ the goal configuration, and $\mathbf{p}$ is the obstacle point in the work space. The proposed cost is defined as:
@@
where $\alpha_{1},\alpha_{2}>0$ are scalar weights. $\theta_{1}\in[0,\pi]$ is the angle between the obstacle normal direction at $\mathbf{q}_{t}$ and the motion vector $\mathbf{q}_{next}=(\mathbf{q}_{t+1}-\mathbf{q}_{t})$, while $\theta_{2}\in[0,\pi]$ is the angle between the goal vector $\mathbf{e}_{t}=(\mathbf{q}_{f}-\mathbf{q}_{t})$ and the motion vector $\mathbf{q}_{next}$. When $\theta_{2}=0$, the robot moves directly toward the goal; $\theta_{2}=\pi$ indicates motion directly away from it. And $\theta_{2}$ can be computed as:
$\theta_{2}(\mathbf{q}_{t},\mathbf{q}_{t+1})$ $=\arccos\frac{\mathbf{q}_{next}\cdot\mathbf{e}_{t}}{\|\mathbf{q}_{next}\|\,\|\mathbf{e}_{t}\|},$ (8)
@@
  • Goal proximity: If the distance to the obstacle is larger than the distance to the goal, i.e., $f_{c}(\mathcal{P},\mathbf{q}_{t})\geq\lVert\mathbf{e}_{t}\rVert$, as illustrated in Fig. 2(Right) and Fig. 3, it is more efficient to move directly towards the goal rather than proactively avoiding the obstacle.
Figure 2: Two-link robot in the configuration space. Irregular shapes represent obstacles, and the black dot denotes the current state of the robot. Left: The blue arrow indicates the actual motion direction of the robot at the next step, forming two angles with the CDF gradient and the goal state. Right: Illustration showing why the obstacle-avoidance angle cost is set to zero when the robot–obstacle distance exceeds the robot–goal distance.
Figure 3: Left: If the robot–obstacle distance exceeds the robot–goal distance while the obstacle cost remains active, the robot may circle around the obstacle. Right: When the obstacle cost is set to zero under the same condition, the robot converges to the goal more directly. The star marker indicates the goal configuration.
This design offers two advantages: (1) all cost terms are expressed in the same unit (angle), making the weight selection more intuitive; (2) obstacle avoidance and goal-seeking behaviors are directly captured in a single unified metric. In MPPI, a batch of $N$ control sequences is sampled from $\pi_{t}$, and each sequence is rolled out over a prediction horizon $H$ to compute the total cost. With the proposed angle-based cost, collision risk and goal progress can be evaluated locally at the current time step, without simulating the full horizon. Therefore, we modify the sampling policy to generate only $N$ single-step control inputs ($H=1$) from $\pi_{t}$, compute $\mathbf{q}_{next}$, and evaluate the cost directly. This reduces computation while retaining effective obstacle avoidance and goal-directed motion.
@@ III-B One-step MPPI for Configuration Space Motion Planning
The projection of any control $\mathbf{u}\in\mathbb{R}^{n}$ onto $\mathcal{U}_{\text{step}}(\mathbf{q})$ is the elementwise clipping
@@ IV Numerical Simulations
In this section, we report simulation results on a two-link robot and a Franka Emika robot to evaluate the proposed method. The experiments are designed to (i) demonstrate the necessity of updating both the mean and covariance of the sampling distribution, (ii) show that naïve sampling without distribution updates is insufficient, (iii) assess the effectiveness of our cost-function design, and (iv) compare our approach against representative baselines. Finally, we provide a qualitative analysis to further validate the method. All experiments were performed on a workstation with an AMD Ryzen 7000 CPU and an NVIDIA RTX 4070 GPU.
@@
0: Initial configuration $\mathbf{q}_{s}$; goal configuration $\mathbf{q}_{f}$; obstacle point cloud $\mathcal{P}$; joint limits $(\mathbf{q}_{\min},\mathbf{q}_{\max})$; time step $\Delta t$; number of samples NN; angle-based cost weights $(\alpha_{1},\alpha_{2})$; obstacle threshold distance $d_{act}$. MPPI parameters.
@@
1: Initialize $\mathbf{q}_{0}\leftarrow\mathbf{q}_{s}$
2: Initialize Gaussian policy $\pi_{0}=\mathcal{N}({\boldsymbol{\mu}}_{0},{\boldsymbol{\Sigma}}_{0})$
@@
4:  Sample NN control inputs $\{\mathbf{u}_{i}\}_{i=1}^{N}\sim\pi_{t}$ {one-step (H=1H=1)}
@@
11:  Update policy parameters ${\boldsymbol{\mu}}_{t+1}$, ${\boldsymbol{\Sigma}}_{t+1}$ using Eq. 1a and Eq. 1b
12:  Set the next control input $\mathbf{u}_{\mathrm{des}}\leftarrow{\boldsymbol{\mu}}_{t+1}$
@@
14:  if $\lVert\mathbf{q}_{t+1}-\mathbf{q}_{f}\rVert<allow\_range$ then
@@
IV-A 2-DoF Two-Link Robot
We first evaluate the proposed method on a 2D two-link robot with link lengths $l_{1}=l_{2}=2$ and joint limits $q\in[-\pi,\pi]$ for each joint. Two circular obstacles are placed at positions (2.3,2.3)(2.3,-2.3) and $(0.0,2.45)$, each with a radius of $0.3$. The start configuration is $\mathbf{q}_{s}=[2.1,1.2]$, and the goal configurations are $\mathbf{q}_{f}=[-2.1,-0.9]$ and [0.5,0.0][-0.5,0.0], and control time step to $\Delta t=0.01s$.
Figure 4: Comparison of different methods in both configuration space and workspace trajectories. The background color encodes the CDF value. Left: Trajectories from the same start configuration but different goals, used to validate the effectiveness of our cost function design. A purely random baseline, which samples controls independently at each step without policy updates, highlights the necessity of distribution updates in MPPI. Right: Comparison with MPPI and an optimization-based method. For QP2 (black), the start and goal are identical to those used in MPPI (Original) and MPPI (Ours), but the trajectory converges to a local optimum and fails to progress. In contrast, QP1, with different start and goal states, successfully reaches the target.
@@
We further provide a comprehensive comparison with other baseline methods: the original MPPI [13] and the QP-based method from [7], solved using IPOPT [14]. The start and goal configurations are the same as in the previous experiment, with an additional test case for the QP method (labeled QP 1), where the start is $[2.5,0.5]$ and the goal is [2.0,0.3][-2.0,0.3]. The cost function used in the 2-DoF MPPI experiment is defined as
$c_{x,u}=\alpha_{g}c_{\text{goal}}+\alpha_{c}c_{\text{col}}+\alpha_{j}c_{\text{joint}}+\alpha_{s}c_{\text{stay}}.$ (16)
where $c_{\text{goal}}$ denotes the terminal cost measuring the error between the predicted final configuration and the target configuration $\mathbf{q}_{f}$, while $c_{\text{stay}}$ penalizes trajectories that remain stationary over time [4].
The trajectories in configuration space are shown in Fig. 4(Right). The QP-based method [7] occasionally fails to reach the goal for certain start–goal pairs, as the control input oscillates between two states, leading to stagnation and convergence to a local optimum (QP 2, with the same start and goal as MPPI (Ours) and MPPI (Original)). For comparison, in QP 1 with a different start–goal pair, the method successfully reaches the goal. The original MPPI produces trajectories that often run close to obstacles and joint limits, and the resulting paths are generally less smooth.
@@
measured in radians in the configuration space. All methods are tested under the same random seed with identical sampling rules for start and goal states, repeated 500 times. The results are summarized in Table. I. Our proposed method achieves the highest success rate. Although it requires more steps and produces slightly longer paths, this can be attributed to its ability to avoid local optima and handle longer trajectories effectively. The failure cases mainly occur when the start configuration is very close to an obstacle, causing collisions before the distribution has been sufficiently updated. In contrast, the original MPPI may still result in collisions with obstacles or violations of joint limits due to the difficulty of appropriately balancing the weights among different cost terms.
@@
$\mathbf{u}=k(\mathbf{q}_{f}-\mathbf{q}_{s}).$ (18)
@@
Figure 6: Comparison between the QP method and MPPI (ours) in Scenario 1. The green marker indicates the start posture, while the red marker indicates the goal posture. Top: Our method successfully avoids the cross obstacle and reaches the goal. Bottom: The QP method becomes trapped near the obstacle, exhibiting oscillatory behavior around it and failing to make further progress due to local optimality.
The results are summarized in Table LABEL:tab:exp2. Our method achieves consistently high success rates. The path length results indicate that, for the QP method, longer trajectories are more likely to become trapped in local optima. An example of the QP method being trapped in a local optimum, leading to oscillatory inputs, is illustrated in Fig. 6. In terms of computational efficiency, our approach achieves an average of 776 Hz in Scenario 1 (with 60 obstacles) and 718 Hz in Scenario 2 (with 123 obstacles).
We further compare the computation frequency with the MPPI method from cuRobo [15]. We do not directly compare our method with the original MPPI in the configuration space, primarily because it is challenging to design a suitable cost function. The original MPPI framework is tailored for end-effector tracking, where multiple joint configurations correspond to the same end-effector pose due to the redundancy of the 7-DOF Franka robot. However, when the terminal cost is defined in the configuration space, the solution becomes unique, which makes it difficult for MPPI to converge. Therefore, we only compare the control frequency in this case. Specifically, we provide MPPI with the end-effector poses corresponding to the target joint configurations as its input. Under this setup, MPPI achieves an average of 61 Hz in Scenario1 and 28 Hz in Scenario2. As the number of obstacles increases, collision-checking time also grows. In contrast, [15] reported that MPPI can reach up to 500 Hz on an RTX 4090 when the number of obstacles is relatively small. Compared to this, our method sustains significantly higher control frequencies while maintaining near-perfect success rates.